OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 72, 18, 9/2 , 9/8 , ... Converges; 0 Converges; -6120 Diverges Converges; 96

terenzreignz (terenzreignz):

First off,what kind of sequence is this?

OpenStudy (anonymous):

I have no clue

terenzreignz (terenzreignz):

This is what's called a "geometric" sequence. What that means is that if you divide the second term by the first, the third term by the second, the fourth term by the third, etc... You should get the same quotient.

OpenStudy (anonymous):

oh ok cool

OpenStudy (tkhunny):

It is impossible to tell convergence without a general term. Four terms does not a definition of in infinite sequence make. The 5th terms may be -2. Where would that leave us?

terenzreignz (terenzreignz):

@tkhunny lol that is true :3 -It is always possible to find a rule which justifies a finite sequence to be continued by any number- --The Oxford Murders But let's stick to the simplest, whatever that means :D

terenzreignz (terenzreignz):

@tholn Divide the second term by the first, what do you get?

OpenStudy (anonymous):

72/18=4

OpenStudy (tkhunny):

Two things: 1) The author of the question should know better, or 2) You are expected to point out this deficiency and then solve the problem with whatever assumptions seem reasonable, such as the fact that the first three ratios are the same.

terenzreignz (terenzreignz):

divide the second term by the first... means second over first @tholn a bit... backwards, lol What I meant would be 18/72 = 1/4 everything all right so far?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay, now divide the third term by the second...

OpenStudy (anonymous):

(9/2)/18 = 1/4

terenzreignz (terenzreignz):

And finally, divide the fourth term by the third (sorry if this is tedious, this'll be the last division, I promise :) )

OpenStudy (anonymous):

(9/8)/(9/2)=1/4

terenzreignz (terenzreignz):

Okay. You'll notice that every time we divided (second over first, third over second, fourth over third, etc) The quotient is 1/4 1/4 is what you call the "common ratio" of the geometric sequence. Now if the common ratio is STRICTLY BETWEEN -1 and 1, then the sequence ALWAYS converges to zero :) IS 1/4 between -1 and 1? \[\Large -1 < \frac14 <1 \qquad\qquad\color{red}{??}\]

OpenStudy (anonymous):

oh ok so it Converges; 0

terenzreignz (terenzreignz):

That is correct ;)

terenzreignz (terenzreignz):

Also note, that if you get a common ratio that it -1 or less , or 1 or more, in other words, if you get a common ratio r such that... \[\Huge |r|\ge 1\] Then the sequence ALWAYS diverges.

OpenStudy (mayaal):

u explain sooo well!